# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 12. Vertex-transitive graphs
[…]
($\Leftarrow$). Clearly, if $G$ is one of the graphs in (i), (iii), (iv), (v), then $G$ is vertex-transitive. If (ii) $G$ is the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph, then $G$ is vertex-transitive by Lemma 12.21. Both (vi) a cycle of $K_{2,2}$-bags and (vii) the line graph of the cube are arc-transitive. If $G$ is a $K_{4}$-expansion of (vi) or (vii), then $G$ is vertex-transitive by Lemma 12.21. ∎
@@ Acknowledgement.
We thank Joseph Devine for exciting discussions. We thank Agelos Georgakopoulos for telling us about Theorem 12.1 of Godsil and Royle.
Appendix A Circular saws
@@
Lemma A.1.
Let $G$ be a $k$-regular circular saw on $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$. Let $aa^{\prime}$ and $bb^{\prime}$ be two edges of $G$ such that $\{a,a^{\prime}\}$ and $\{b,b^{\prime}\}$ are disjoint and there is no $\{a,a^{\prime}\}$$\{b,b^{\prime}\}$ edge. Let $X$ denote the neighbourhood of $\{a,a^{\prime}\}$, and let $Y$ denote the neighbourhood of $\{b,b^{\prime}\}$. There are $2(k-1)$ disjoint $X$$Y$ paths in $G$.
@@
For each vertex $x\in X$ we construct two $x$$Y$ paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See Figure A.1 for an example.
@@
Exactly one of the paths $P^{+}_{x},P^{-}_{x}$ has an internal vertex in $X$. Let $P_{x}$ denote the other path. Then $(P_{x}:x\in X)$ is a family of disjoint $X$$Y$ paths. There are $|X|=2(k-1)$ of these paths. ∎
Figure A.1. Paths constructed in the proof of Lemma A.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
Lemma A.2.
@@
We show first that $G$ is $k$-connected, and assume for a contradiction that some two vertices $a,b$ of $G$ are separated by a set $S$ of $k-1$ vertices. Let $a^{\prime}$ be a neighbour of $a$ outside $S$, and let $b^{\prime}$ be a neighbour of $b$ outside $S$. Note that $12(k-1)\geqslant k+$ 1 since $k\geqslant 3$. Hence by Lemma A.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$. Hence $G$ is $k$-connected.
Next, we assume for a contradiction that $G$ has a $k$-separation $(A,B)$ such that $S(A,B)$ fails to be the neighbourhood of some vertex of $G$. Let $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Then there are vertices $a^{\prime}\in N(a)\smallsetminus S(A,B)$ and $b^{\prime}\in N(b)\smallsetminus S(A,B)$. Now Lemma A.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$. ∎
Appendix B Comparison of strict tri-separations and tri-separations
The strict tri-separation-decomposition (2) is different from the tri-separation decomposition [Tridecomp, Theorem 1] as can be seen in Figures B.1 and B.2. However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via Lemma B.1.
Figure B.1. The totally-nested strict tri-separations
Figure B.2. The totally-nested non-trivial tri-separations
@@ Appendix B Comparison of strict tri-separations and tri-separations
Lemma B.1.
Let $G$ be a 3-connected graph. Let $(A,B)$ be a totally-nested non-trivial tri-separation of $G$. Then the left-right-shift $(A^{\prime},B^{\prime})$ of $(A,B)$ is a totally-nested strict tri-separation, or $(A,B)$ is negligible.
@@
Assume that $(A,B)$ is not negligible. Then $(A^{\prime},B^{\prime})$ is a strict tri-separation by 5.10 (which generalises to strict tri-separations via an apex-argument as in the proof of Corollary 3.2). Assume for a contradiction that there is a strict tri-separation $(C,D)$ of $G$ that crosses $(A^{\prime},B^{\prime})$ with $C$-link $\gamma^{\prime}$ and $D$-link $\delta^{\prime}$. We will show that $(C,D)$ crosses $(A,B)$ with non-empty $C$-link $\gamma$ and non-empty $D$-link $\delta$. For this, we consider the crossing-diagram of $(C,D)$ and $(A^{\prime},B^{\prime})$. By [Tridecomp, Lemma 1.3.10], there are two cases.
@@
Case: the centre consists of three vertices and all links are empty. Then $S(A^{\prime},B^{\prime})$ consists of three vertices, so $(A^{\prime},B^{\prime})=(A,B)$. But all corners are non-empty, so $(A^{\prime},B^{\prime})$ is not half-connected. However, $(A,B)$ is half-connected because it is totally-nested, by [Tridecomp, Lemma 2.3.4], contradicting that $(A^{\prime},B^{\prime})=(A,B)$. ∎
[…]

### v2

@@ 13. Vertex-transitive graphs
Now assume that $H^{\prime}_{t}$ is bad. Then $V(H^{\prime}_{t})$ consists of exactly five vertices. Let $v$ be the vertex in the interior of $H^{\prime}_{t}$. By 4-regularity of $L$, the vertex $v$ is adjacent to every vertex in $H^{\prime}_{t}-v$. Since every torso of $\mathcal{O}^{\prime}$ is 3-connected, every vertex $u$ in an adhesion-set of $H^{\prime}_{t}$ has degree 2 in $G^{\prime}_{t}$. It follows that the torso $H^{\prime}_{t}$ is a 4-wheel, that $G^{\prime}_{t}$ is the union of two triangles that meet exactly in $v$, and that $V_{t}$ induces a $K_{2}$ in $G$.
(iii). We denote the vertices in $U_{i}$ by $u_{i}$ and $u_{i}^{\prime}$. Assume for contradiction that $V_{t}$ induces a 4-cycle $Z$ that does not alternate between $U_{1}$ and $U_{2}$. As shown in the proof of (ii), the tetra-cut $(V_{t},V(G)\smallsetminus V_{t})$ is totally-nested. Then without loss of generality $Z=u_{1}u_{1}^{\prime}u_{2}^{\prime}u_{2}u_{1}$. Then, however, $(V_{t_{1}}\cup U_{1},V(G)\smallsetminus(V_{t_{1}}\cup U_{1}))$ is a tetra-cut that crosses $(V_{t},V(G)\smallsetminus V_{t})$, a contradiction. ∎
Key Lemma 13.44.
Let $G$ be a 3-regular internally-4-connected vertex-transitive graph. Assume that $G$ has two crossing tetra-cuts. Then $G$ is a cycle of $C_{4}$-bags of length $\geqslant 4$, or a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 8$.
Proof.
By Lemma˜13.43GG admits a cycle-partition $\mathcal{O}=(O,(V_{t})_{t\in O})$ with $|O|\geqslant 4$ such that properties (i)(iii) from Lemma˜13.43 are fulfilled. Let $O=:t_{1}t_{2}\ldots t_{\ell}t_{1}$. For $t=t_{i}\in V(O)$ we denote by $U_{t,1}$ the set of endvertices of the edges in the adhesion-set for $tt_{i-1}$ that lie in $V_{t}$; and by $U_{t,2}$ we denote the set of endvertices of the edges in the adhesion-set for $tt_{i+1}$ that lie in $V_{t}$.
We show that either every bag $V_{t}$ of $\mathcal{O}$ induces a $K_{2}$ or every bag $V_{t}$ of $\mathcal{O}$ induces a $C_{4}$ alternating between $U_{t,1}$ and $U_{t,2}$. In the former case, it follows that $G$ is a cycle of $C_{4}$-bags of length $\geqslant 4$, while in the latter case it follows that $G$ is a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 8$.
Assume for a contradiction that there is a bag of $\mathcal{O}$ that induces a $K_{2}$, and that there is a bag of $\mathcal{O}$ that induces a $C_{4}$. Then, without loss of generality, one of the following cases occurs.
The bags for $t_{1}$, $t_{2}$ and $t_{3}$ induce a $K_{2}$ while the bag for $t_{4}$ induces a $C_{4}$. Then every vertex in $V_{t_{2}}$ is contained in two $C_{4}$’s of $G$ while every vertex in $V_{t_{4}}$ is contained in exactly one $C_{4}$ of $G$, contradicting vertex-transitivity.
The bags for $t_{2}$ and $t_{3}$ induce a $K_{2}$ while the bags for $t_{1}$ and $t_{4}$ induce a $C_{4}$. Let $u\in V_{t_{2}}$. Then $u$ is contained in exactly two 5-cycles $Z,Z^{\prime}$. The intersection $Z\cap Z^{\prime}$ is a path with three edges, and $u$ is an internal vertex of this path. The neighbour of $u$ in $V_{t_{1}}$ also lies on $Z$ and $Z^{\prime}$, but it is not an internal vertex of $Z\cap Z^{\prime}$, contradicting vertex-transitivity.
The bag for $t_{2}$ induces a $K_{2}$ while the bags for $t_{1}$ and $t_{3}$ induce a $C_{4}$. Then every vertex in $V_{t_{2}}$ is contained in exactly four 5-cycles of $G$, while every vertex in $V_{t_{1}}\cap N(V_{t_{2}})$ is contained in at most three 5-cycles. This contradicts vertex-transitivity. ∎
Recall that a graph $G$ is 2-quasi-5-connected if $G$ is quasi-4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|\leqslant 2$ or $|B\smallsetminus A|\leqslant 2$.
Lemma 13.45.
Let $G$ be a 3-regular quasi-4-connected graph. Assume that $G$ has no tetra-cut. Then $G$ is 2-quasi-5-connected.
Proof.
If $G$ is not 2-quasi-5-connected, it has a 4-separation $(A,B)$ with $|A\smallsetminus B|\geqslant 3$ and $|B\smallsetminus A|\geqslant 3$. Since $G$ is quasi-4-connected, it follows that $(A,B)$ is tight. Let $S(A,B)=\{v_{1},v_{2},v_{3},v_{4}\}$.
First we show that if $v_{i}$ only has one neighbour $u_{i}$ in $A\smallsetminus B$ and $v_{j}\neq v_{i}$ only has one neighbour $u_{j}$ in $A\smallsetminus B$, then $u_{i}\neq u_{j}$. Indeed, otherwise $(A^{\prime},B^{\prime}):=((A\cup\{u_{i}\})\smallsetminus\{v_{i},v_{j}\},B\cup\{u_{i}\})$ is a 3-separation of $G$ with $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$, contradicting that $G$ is quasi-4-connected. By the above argument, it follows that the left-right-reduction of $(A,B)$ satisfies the matching-condition.
Since $G$ is 3-regular and $(A,B)$ is tight, every vertex $v_{i}\in S(A,B)$ either has exactly one neighbour in $A\smallsetminus B$, or $v_{i}$ has exactly one neighbour in $B\smallsetminus A$ and no neighbour in $A\cap B$. Therefore, the left-right-reduction of $(A,B)$ is a 4-edge-cut and hence a tetra-cut of $G$. ∎
Example 13.46.
The toroidal hex-grid on a torus shows that 2-quasi-5-connected in Lemma˜13.45 is best possible.
Lemma 13.47.
Let $G$ be a 3-regular internally-4-connected vertex-transitive graph such that every tetra-cut of $G$ is totally-nested. Then either $G$ is 2-quasi-5-connected or there is a canonical partition $\mathcal{P}=\{\,A_{i}:i\in I\,\}$ of $V(G)$ such that
  1. (i)
    each $A_{i}$ induces a $C_{4}$;
  2. (ii)
    the edges in $G$ with ends in different classes of $\mathcal{P}$ form a perfect matching in $G$.
Proof.
If $G$ has no tetra-cut, then $G$ is 2-quasi-5-connected by Lemma˜13.45. So we may assume that $G$ has at least one tetra-cut. Let $\sigma:=\{\,(A_{i},B_{i}):i\in I\,\}$ be the set of all tetra-cuts of $G$ whose side $A_{i}$ induces a 4-cycle. Since all tetra-cuts of $G$ are totally-nested, $\sigma$ is a star of tetra-cuts. Hence the sets $A_{i}$ are pairwise disjoint. Since $G$ is vertex-transitive, we have $\bigcup_{i\in I}A_{i}=V(G)$. Therefore, $\mathcal{P}:=\{\,A_{i}:i\in I\,\}$ is a canonical partition of $G$ into vertex-sets that induce 4-cycles.
Let $M$ denote the set of all edges in $G$ with ends in different classes of $\mathcal{P}$. Every vertex $v$ of $G$ lies in exactly one $A_{i}$, and since $G$ is 3-regular, it follows that $v$ is incident to exactly one edge in $M$. Hence $M$ is a perfect matching. ∎
Key Lemma 13.48.
Let $G$ be a 3-regular internally-4-connected vertex-transitive graph. Assume that every tetra-cut in $G$ is totally-nested. Then $G$ is one of the following:
  • $\cdot$
    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length 6;
  • $\cdot$
    one of the three $C_{4}$-expansions of the line graph of the cube shown in Figure˜1.15;
  • $\cdot$
    a vertex-transitive $C_{4}$-expansion of a cycle of $K_{2,2}$-bags of length $\geqslant 3$;
  • $\cdot$
    a $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph;
  • $\cdot$
    2-quasi-5-connected.
Proof.
Assume that $G$ is not 2-quasi-5-connected. Then $V(G)$ has a canonical partition $\mathcal{P}$ into vertex-sets that induce 4-cycles by Lemma˜13.47. Fix $U\in\mathcal{P}$ and let $\{\,U_{i}:i\in I\,\}$ be the set of classes $U_{i}\in\mathcal{P}$ that contain neighbours of $U$. Let $M$ denote the set of edges of $G$ with ends in different classes of $\mathcal{P}$, so $M$ is a perfect matching by Lemma˜13.47. Let $F\subseteq M$ denote the set of edges leaving $U$.
Case: all four edges in $F$ end in the same $U_{i}$. Then $F=M$ and so $V(G)=U\cup U_{i}$. Both $U$ and $U_{i}$ induce 4-cycles, so either $G=C_{4}\square K_{2}$ which is a cycle of $C_{4}$-bags of length 4, or $G$ is the life belt. However, in both cases $G$ has two crossing tetra-cuts, so neither case occurs.
Case: exactly three edges in $F$ end in the same $U_{i}$. This case is impossible: Then the vertex in $U$ and the vertex in $U_{i}$ that are not incident with these three edges form a 2-separator of $G$, contradicting that $G$ is internally-4-connected.
Case: there is a $U_{i}$ in which exactly two edges from $F$ end. Then four edges from $M$ leave $U\cup U_{i}$. Hence $U\cup U_{i}$ is a side of a tetra-cut of $G$. Since this tetra-cut is totally-nested by assumption, it has a side that induces a 4-cycle by Lemma˜13.38. This side can only be $V(G)\smallsetminus(U\cup U_{i})$, which then must be a $U_{j}$ for some $j\in I\smallsetminus\{i\}$. Now $U$ and $U_{i}$ each send exactly two edges to $U_{j}$. If $G$ is a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length 6, then we are done. Otherwise the 4-cycle $O$ induced by $U$, say, first traverses the two neighbours $u_{i},u_{i}^{\prime}$ of $U_{i}$ in $U$, and then traverses the two neighbours $u_{j},u_{j}^{\prime}$ of $U_{j}$ in $U$. But then (U,V(GU))(U,V(G-U)) and $(U_{i}\cup\{u_{i},u_{i}^{\prime}\},U_{j}\cup\{u_{j},u_{j}^{\prime}\})$ are two crossing tetra-cuts in $G$, contradicting our assumption.
Case: $|I|=4$ and exactly one edge from $F$ ends in each $U_{i}$. Then $G/\mathcal{P}$ is 4-regular and 4-connected. Moreover, $G/\mathcal{P}$ is arc-transitive, as $\mathcal{P}$ is unique and every vertex of $G$ is incident with a unique edge in $F$. Then by Corollaries˜13.6 and 13.25, the graph $G/\mathcal{P}$ is one of the following:
  • $\cdot$
    the line-graph of the cube,
  • $\cdot$
    a cycle of $K_{2,2}$-bags of length $\geqslant 3$, or
  • $\cdot$
    essentially-5-connected.
Assume first that $G/\mathcal{P}$ is the line-graph of the cube, so $G$ is some $C_{4}$-expansion thereof. Since $G$ is vertex-transitive, there are only three possible $C_{4}$-expansions: these are precisely the three depicted in Figure˜1.15.
Assume finally that $G/\mathcal{P}$ is essentially-5-connected. Then $G/\mathcal{P}$ is quasi-5-connected by Corollary˜13.28. Hence $G$ is the $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph. ∎
Proposition 13.49.
A 3-regular quasi-4-connected finite graph $G$ is vertex-transitive if and only if $G$ is either
  1. (i)
    2-quasi-5-connected and vertex-transitive,
  2. (ii)
    a vertex-transitive $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph,
  3. (iii)
    a cycle of $C_{4}$-bags of length $\geqslant 4$ or $K_{3}\square K_{2}$,
  4. (iv)
    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 6$,
  5. (v)
    one of the three $C_{4}$-expansions of the line-graph of the cube shown in Figure˜1.15, or
  6. (vi)
    a vertex-transitive $C_{4}$-expansion of a cycle of $K_{2,2}$-bags of length $\geqslant 3$.
Proof.
($\Rightarrow$). By ˜13.30, we may assume that $G$ is internally 4-connected and has girth $\geqslant 4$. If $G$ has two crossing tetra-cuts, then we are done by ˜13.44. Otherwise every tetra-cut of $G$ is totally-nested. Then we are done by ˜13.48.
($\Leftarrow$) is straightforward. ∎
Proof of Figure˜1.8.
Proposition 13.50.
[CK23, Corollary 2]. Every vertex-transitive finite connected $G$ is either
  • $\cdot$
    3-connected and every non-trivial strong tri-separation of $G$ is a 3-edge-cut with a side that induces a triangle,
  • $\cdot$
    a cycle, $K_{2}$ or $K_{1}$.
Proposition 13.51.
Every vertex-transitive finite connected graph $G$ is either
  • $\cdot$
    internally 4-connected and vertex-transitive,
  • $\cdot$
    the $K_{3}$-expansion of a quasi-4-connected 4-regular arc-transitive graph,
  • $\cdot$
    a cycle, $K_{3}\square K_{2}$, $K_{2}$ or $K_{1}$.
Proof.
By Proposition˜13.50, we may assume that $G$ is 3-connected and every non-trivial strong tri-separation of $G$ is a 3-edge-cut with a side that induces a triangle. If $G$ has no non-trivial strong tri-separation, then $G$ is internally-4-connected by [CK23, 1.2.8]. Hence we may assume that $G$ has at least one non-trivial strong tri-separation, and by assumption it has a side that induces a triangle.
Let $\{\,(A_{i},B_{i}):i\in I\,\}=\sigma$ be the set of all non-trivial strong tri-separations of $G$ with $G[A_{i}]=K_{3}$. Since each $S(A_{i},B_{i})$ is a 3-edge-cut, the tri-separations $(A_{i},B_{i})$ are totally-nested by [CK23, 1.3.10], so $\sigma$ is a star of tri-separations. Hence the $A_{i}$ are pairwise disjoint, and every vertex of $G$ is contained in some $A_{i}$ by vertex-transitivity, so $\mathcal{P}:=\{\,A_{i}:i\in I\,\}$ is a partition of $V(G)$.
Since each $S(A_{i},B_{i})$ is a 3-matching, the set $M$ of cross-edges of $\mathcal{P}$ is a perfect matching. If two edges in $M$ join $A_{i}$ to $A_{j}$ for some $i\neq j\in I$, then the two vertices in $A_{i}\cup A_{j}$ that are not incident to these edges form a 2-separator of $G$ contradicting 3-connectivity, unless $G=K_{3}\square K_{2}$. Hence we may assume that at most one edge in $M$ joins any $A_{i}$ to any $A_{j}$, so $G/\mathcal{P}$ is 3-regular.
Since $G$ is vertex-transitive and the $K_{3}$-expansion of $G/\mathcal{P}$, it follows that $G/\mathcal{P}$ is arc-transitive by Lemma˜13.21. In particular, $G/\mathcal{P}$ is vertex-transitive, and since $G/\mathcal{P}$ is 3-regular we get that $G/\mathcal{P}$ is 3-connected by Proposition˜13.50. However, $G/\mathcal{P}$ has no non-atomic 3-edge-cut, as this would yield a non-trivial strong tri-separation of $G$ that is a 3-edge-cut but whose sides have size $\geqslant 9$, contradicting our assumption. Hence applying Proposition˜13.50 to $G/\mathcal{P}$ in combination with [CK23, 1.2.8] yields that $G/\mathcal{P}$ is quasi-4-connected. ∎
@@ Acknowledgement.
We thank a referee for valuable comments on the introduction. We thank another referee for bringing [LPPS22] to our attention. We thank Joseph Devine for exciting discussions. We thank Agelos Georgakopoulos for telling us about Theorem˜13.1 of Godsil and Royle.
14. Appendix A: Circular saws
@@
Lemma 14.1.
Let $G$ be a $k$-regular circular saw on $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$. Let $aa^{\prime}$ and $bb^{\prime}$ be two edges of $G$ such that $\{a,a^{\prime}\}$ and $\{b,b^{\prime}\}$ are disjoint and there is no $\{a,a^{\prime}\}$$\{b,b^{\prime}\}$ edge. Let $X$ denote the neighbourhood of $\{a,a^{\prime}\}$, and let $Y$ denote the neighbourhood of $\{b,b^{\prime}\}$. There are 2(k1)2(k-1) disjoint XXYY paths in $G$.
@@
For each vertex $x\in X$ we construct two xxYY paths $P^{+}_{x},P^{-}_{x}$ inductively, as follows. See Figure˜14.1 for an example.
@@
Exactly one of the paths $P^{+}_{x},P^{-}_{x}$ has an internal vertex in $X$. Let $P_{x}$ denote the other path. Then $(P_{x}:x\in X)$ is a family of disjoint XXYY paths. There are |X|=2(k1)|X|=2(k-1) of these paths. ∎
Figure 14.1. Paths constructed in the proof of Lemma˜14.1. The edges $aa^{\prime}$ and $bb^{\prime}$ are coloured blue, the paths are coloured red.
Lemma 14.2.
@@
We show first that $G$ is $k$-connected, and assume for a contradiction that some two vertices $a,b$ of $G$ are separated by a set $S$ of k1k-1 vertices. Let $a^{\prime}$ be a neighbour of $a$ outside $S$, and let $b^{\prime}$ be a neighbour of $b$ outside $S$. Note that $2(k-1)\geqslant k+1$ since $k\geqslant 3$. Hence by Lemma˜14.1, there are $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $|S|\leqslant k-1$. Hence $G$ is $k$-connected.
Next, we assume for a contradiction that $G$ has a $k$-separation $(A,B)$ such that $S(A,B)$ fails to be the neighbourhood of some vertex of $G$. Let $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Then there are vertices $a^{\prime}\in N(a)\smallsetminus S(A,B)$ and $b^{\prime}\in N(b)\smallsetminus S(A,B)$. Now Lemma˜14.1 yields $k+1$ disjoint $N(\{a,a^{\prime}\})$$N(\{b,b^{\prime}\})$ paths in $G$, contradicting that $(A,B)$ has order $k$. ∎
15. Appendix B: Comparison of strict tri-separations and tri-separations
The strict tri-separation-decomposition (˜2) is different from the tri-separation decomposition [CK23, Theorem 1] as can be seen in Figures˜15.1 and 15.2. However, the strict tri-separation-decomposition can be viewed as a refinement of the tri-separation decomposition via Lemma˜15.1.
Figure 15.1. The totally-nested strict tri-separations
Figure 15.2. The totally-nested non-trivial tri-separations
@@ 15. Appendix B: Comparison of strict tri-separations and tri-separations
Lemma 15.1.
Let $G$ be a 3-connected graph. Let $(A,B)$ be a totally-nested non-trivial tri-separation of $G$. Then the left-right-reduction $(A^{\prime},B^{\prime})$ of $(A,B)$ is a totally-nested strict tri-separation, or $(A,B)$ is negligible.
@@
Assume that $(A,B)$ is not negligible. Then $(A^{\prime},B^{\prime})$ is a strict tri-separation by ˜5.10 (which generalises to strict tri-separations via an apex-argument as in the proof of Corollary˜3.2). Assume for a contradiction that there is a strict tri-separation $(C,D)$ of $G$ that crosses $(A^{\prime},B^{\prime})$ with $C$-link $\gamma^{\prime}$ and $D$-link $\delta^{\prime}$. We will show that $(C,D)$ crosses $(A,B)$ with non-empty $C$-link $\gamma$ and non-empty $D$-link $\delta$. For this, we consider the crossing-diagram of $(C,D)$ and $(A^{\prime},B^{\prime})$. By [CK23, Lemma 1.3.10], there are two cases.
@@
Case: the centre consists of three vertices and all links are empty. Then $S(A^{\prime},B^{\prime})$ consists of three vertices, so $(A^{\prime},B^{\prime})=(A,B)$. But all corners are non-empty, so $(A^{\prime},B^{\prime})$ is not half-connected. However, $(A,B)$ is half-connected because it is totally-nested, by [CK23, Lemma 2.3.4], contradicting that $(A^{\prime},B^{\prime})=(A,B)$. ∎